Find using distributive property: 
728 × 101 ?????????????

Grade:12

1 Answers

Harshit Singh
3 years ago
Dear Student

Given 728 × 101
=728 × (100 + 1)
= 728 × 100 + 728 × 1
= 72800 + 728
= 73528

Thanks